System modification history
This page records the dates when
modifications were made to the system that may affect observations.
It is not necessarily complete and the dates may be only
approximate. For a history of turret positions see turret
Bottom of page (most recent entry).
- 21sep98 hurricane georges.
- jan99 Dome lowered about 3 ".
- 01feb00 elevation
rails reshimmed ao9 survey.
- 11feb00 Pointing model 11 completed and
- 11feb00 turret positions of horns resurveyed
by lynn baker. Turret positions updated.
- 07mar00 Started tracking the average height
of the platform using the distomats and tiedowns.
- 04sep00 pointing preprocessor installed and
- 23aug00 610 receiver model was updated.
Had been using lbn model. Checked
the pointing and it was off by 5 arcminutes (5db loss). see
610 calibration to see how good the model is now.
- xxnov00 430 gregorian dome switched from
linear to circular polarization.
- xxdec00 Wheels on dome broke. Replaced them.
Support beam (2000 lbs) now on az arm.
- 13jan01 Reflector tieback adjustments below
main support cables completed (1st cycle)
2000 points adjusted. This is the
- 16mar01 Completed pointing model 12 (but not installed).
Pointing model 12 taken during the
day with active tracking of average platform height.
Verification of the model didn't look so good. They were
some unexplained jumps in the pointing that didn't fit well.
Sband used this model for a few observations.
- mar01 ?? Reflector tieback adjustments below main support
cables completed (2nd cycle).
The date is a little suspect. For
sure it is before 20 apr01 since the remeasurements were
done by then.
- 23may01 610 feed changed from skinny to fat dipole feed. (day
- 01jul01 Completed reflector panel adjustments (all
panels) SE quadrant of dish.
- 25jul01 Completed reflector panel adjustments (all
panels) NW quadrant of dish (date??).
- 05jul01 610 model of 23aug00 was not being used by
the gui. Gui updated to use model.
The gui was still using the old lbn
model. The model from 23aug00 was not being used so users of
610 system via the gui were still 5 arcminutes off (5db
down). Updated the gui to use the correct model today.
- 31jul01 tertiary motors installed.
- 01aug01 theodolite survey. Tertiary and dome paddle
- 04aug01 tilt sensor run all az,za to meausure
- 09aug01 theodolite survey of dome from ao9.
- 10aug01 angle of teflon pads on tertiary adjusted so
pads are parallel to rubbing blocks.
- 15aug01 panel adjustments for northwest quadrant of dish
- 19aug01 14:00 ast. updated dut1 (utc-ut1). Was off
by -.125 seconds of time .
- 01sep01 updated sbw hcorcal on 04apr02 retroactive
to this date.
- 16nov01 lband wide filter bank filter #6 (1360-1470)
were removed. The filters were bad (not sure for how long they
had been bad.
- 15dec01 sbn. updated gain(az,za) in idl/data for
gainget() routine. uses sep01->nov01 observations after dish
- 19dec01 xband receiver installed.
- 27dec01 heating resistor installed on xband waveguide
- 06feb02 xband horn moved up .44 inches.
turret position 161.99->162.34, model updated 15.asecs (from
horn survey 05feb02). Then switch xband model to model 13a
- 27feb02 dome weighed using load cells and
- 04mar02 bill sisk adjusted mindex on the
tertiary. It has moved to the edge of the window.
- 20mar02 switch new ac in dome fixed.Both ac
now on. old:19C, new 22 C
- 29mar02 surveyed sbh,xb,sbn.. sbh needs to move 1.068 in y,
xband -1.131 in y
- 31mar02 surveyed the tertiary. z error -.2747 inches.
- 31ma02 sbh moved turret position 125 to 124.150
from horn survey (sidelobes then looked bad).
- 02apr02 xband. moved turret position
162.34->161.50 and updated model offset. From 29mar02 survey
- 02apr02 sbh. turret pos 124.15->125.75 (from
- 03apr02 sbh. turret pos 125.75->125.827 got
survey sign correct. SBH az offset +21 -3.45 from 02apr02 pnt
check and move of turret position.
- 04apr02 sbw. added new calvalue for hcorcal.
Data from sep01. backdata active date to 01sep01. The diodes had
been replaced prior to this.
- 09jul02 15:00 hydraulic brake on dome now
- 20jul02 327 installed new rcvr
..room temp amps...
- 01aug02 around jul02,aug02 felipe tightend
the small rollers for the dome on the right side. Play is now
about 1/32 inch.
- 15aug02 327 rcvr pointing model updated
offset az,za +3,+130" fromcalib
- 15aug02 610 rcvr. filter retuned to 609.5
Mhz 4Mhz Bw. updated calib config, tsysconfig, ifsetup.proc.
- 21aug02 lbn horn focus moved down 3.84
inches to correct for wrong
- 28aug02 cb horn focus moved down 1.08 inches
to correct for wrong
- 30aug02 610 antenna switched from fat
dipolse back to skinny dipole.
- 04sep02 outside rollers on elevation rail
for dome moved in 1/4" to tighten up play as dome moves
- 05sep02 sband high shutter installed.
- 10sep02 new lbn gain curves using data
21aug02->9sep02. Za only fit. Good from 21aug02 when
horn was moved.
- 15sep02 az encoder on ch side broke, was
- 16sep02 az encoder dome side jumped (maybe
when ch side broke). added -1446 to sinza az term for all models
except ch (see 25sep02)
- 17sep02 sbn receiver down for repairs.
lbn,xb dewar temps decrease (sbn is on same compressor).
- 18sep02 1st new whiffltree installed on
dome. Looking uphill it was downhill, right side.
- 21sep02 updated dut1. went from -302 Ms to
-232 Ms (correct) so pointing changed by 70 ms.
- 25sep02 updated carriage house model for az
encoder jump that occured on 16sep02.
- 26sep02 upstairs pwrmeter broke.
swapped upstairs,downstairs pwrmeters.
downstairs using just polB to adjustlevels till new meter
shows up (probably) on 30sep02).
- 04nov02 left downhill dome trolley replaced
with new one.
- 19nov02 new dome trolley installed. looking
uphill 2nd on the right from the bottom
- 20nov02 new dome trolley installed. looking
uphill 2nd on the left from the bottom
- 21nov02 refocused lband wide feed.
- 25nov02 100 Mhz for system found unlocked at
17:45 .This locks everything upstairs. Was jumping by about 200
hz at 12Ghz every 15 seconds. Seti people noticed it
during a test. The pll module that converted from 5 Mhz maser to
100Mhz was out of lock. Fixed 28nov02 17:00. Was locked on
22nov02 for the vlibi run. After that we're not sure.
- 9dec02 5th trolley installed (need to check
- 16dec02 5th trolley removed for adjustments.
- 24dec02 5th trolley reinstalled. (3rd
- 08jan03 6th trolley installed right 4th
uphill. 123 left,124 right
- 15jan03 7th trolley installed right
uphill 3rd position. 123 left, 1234 right.
- 06feb03 lbw receiver removed from turret
floor (placed on tertiary platform) about 400 lbs.
- 08feb03 felipe finished first shimming of
the elevation rails.
- 09feb03 tried to straighten the azimuth arm
"bend" by pulling on the arm....(bend was not really there).
- 10feb03 pnt vme crate power supply expires.
on replacement fiber for irig to platform broke. pnt started
running on its internal clock (on the irig board). By 16feb03 it
had drifted off by about 1.5 seconds.
- 12feb03 survey dome from ao9. 2 za strips, 1
- 13feb03 felipe "touched up" the shims on the
elevation rails during the day.
- 15feb03 more shimming of elevation rails
- 16feb03 18:00. irig fiber to platform fixed.
Pnting was running on local clock. It was 1.5 seconds ahead when
it was fixed. It was probably broken 10feb03 when the pnt vme
crate power supply was changed and the fiber got bumped.
- 17feb03 lbw dewar (about 400lbs) on tertiary
platform lowered to ground. New lbw receiver lifted and
installed in rotary floor
- 21feb03 zeeman box removed from iflo
- 25feb03 shimming of elevation rails
completed. tugging of shim pack horizontally also done.
- 28feb03 added weight (water) to platform.
T12:2400,T4:2400, T8:1400 to compensate for T8 hairpin
- 27feb03 dome lifted about 1.65 inches.
- 01mar03 model14 started. They still have
dome bolts to replace on the dome side. but shimming and dome
motion is done.
- 04mar03 model14 installed with old feed
- 07mar03 18:00 new offsets for model 14
- 10mar03 17:00. offsets of 07mar03 were
entered backwards.. corrected here.
- 31mar03 lbw new calvalues installed
(measured 06mar03, back dated to 17feb03).
- 22apr03 new cal diodes installed on cband
- 23apr03 downstairs iflo polA sbc1 monitor
port isolation improved with another amplifier.
- 05may03 seti IF cabinet removed from turret
floor. 450 lbs.
- 05may03 loose cable in cband cal box
tightened. (or maybe it was 06may03)
- 13may03 lbw. switched order of postamps.
before: 24dbGain,filterBand,10dbGain. after 10dbGain,filterbank,
24dbgain. thought 24dbgain amp causing long compression.
- xxmay03 1320Mhz hipass filter lbw moved from
outside filter bank to inside filter bank. Date is aprrox. The
filter was installed with the polarizations reversed. It was
- 14may03 installed remeasured calvalues for sbh. back
dated to 01mar03.
- 16may03 measured cal values cband using sky and absorber
(cal cable tightened on 6may03)
- 18may03 installed cband cal values 22apr03 to 4may03 and
05may03 to ..
- 30may03 recomputed xband cals using correct definition
- 01jun03 3 filters in cor chassis
replaced/repaired: 1b,4a different filters, 4b repaired filter
from 1b. Eddy ordered 8 new filters to replace them (they were
shorting out and causing the instabilities that people had been
seeing in the correlator bandpasses for the last few months..
- 09jul03: ao9 survey of dome.
- 11jun03 (check date).. bill replaced power
supply for power counters in interim correlator.
- 14jun03 idl lbw cals. date reversed in file
(48 2003 instead of 2003 48). people using idl
31mar03->13jun03 would get the latest cals even if there data
was before this (since year 48 is alway before the current
date). Fixed on this day.
- 14jun03 lbw gain curve. The gain curve of
nov21 2002 after horn moved did not have the date field
(!2003 dayno). People were getting the gain curve from before
the horn move. The file had been edited on 17feb03 so it may
have been removed at that time. Fixed the problem today.
- 18jun03 wapp pulsar observing . polarization
mode , 3level, bw<=50Mhz bug was fixed today. Prior to this,
only polA crosscorrelations were being recorded.
- 26jun03 interim correlator. All 8 lowpass
filters (before digitizers) were replaced with new filters (see
- 27jun03 interim correlator. amp 4b in
250->baseband chassis replaced. It was oscillating.
- 16jul03 ri digitizer Q1 was occasionally
losing the rdFifo pulse. The pulse len was lengthed on this date
and the problem appears to have gone away. This problem was seen
in data as far back as feb02.
- 20jul03 installed new
xband cal values. measured using sky and absorber.
backdated to 01jan03.
- 21jul03 xb updated
pointing offsets from calibration runs.
- 28jul03 lightning
tiedowns. td8: dac blown out. Drove all the way up, limit
switch did not turn it off. replaced dac board, checked limit
switch. Td4: encoder board bad reading 24 inches. replaced
board. Position of td moved by about .8 inches from when it went
off. Someone must have moved it when board was off. Since abs
encoder ok, position is probably ok.
- 04aug03 updated dut1 for pointing. was about 100 millsecs off.
- 05aug03 05,06,07.. moved all the limit
switches on tiedowns. (more
- 15aug03 (approximate date). re enforcement
for t8 corner where main cable cut was installed. started
mods. Added new stuff, removed lead to keep weight the same
- 28aug03 vlba filter 1250-1750 in downstairs
- 3aug03 xband default mixer switched from
10ghz upconverter to 2_12ghz mixer
- 30aug03 polA downstairs iflo 5volts was
shorting out when monitor port sbc1 was touched. The semi rigid
cable inside box had fallen down and was shorting against vcc of
one of the switches inside. This line had been installed
23apr03. This problem may have been occurring for awhile.
- 30apr03 power dip 8am. looks like 11.5 Ghz
lo used to downconvert from 10ghz to 1.5 ghz lost lock. Problem
remained for 23apr03, 24apr03 daytime.
- 03sep03 last part of T8 reinforcement added
335 lbs extra in T8 corner.
- 04sep03 new xband horn installed.
- 04sep03 xband cal cables "cleaned". cal
values no longer valid. need to be remeasured.
- 05oct03 xband rcvr up to test range, cals
tsys measured on sky/load
- 10oct03 installed new sbw
cal values. backdated to 15aug03.
- 14oct03 8:02 am gregAz encoder jumped by 3
teeth (1.66Amin little circle). Someone probably bumped it.
Fixed the problem 16oct03 14:30.
- 14oct03 sbn turnstile stubs switched. Also
reversed the rf cables from dewar. The oc pol now uses the
better amplifier. cals values in tables also switched.
- 27oct03 new xband cal values installed. back
dated to 05oct03
- dec03 Isolation of if2 buffer
amplifier outputs increased via more amps/pads. Now the same as
- 07jan04 lbn receiver removed from dome.
- 07jan04 327 recevier moved from turret=21.5
to turret (about) 342.5
- 08jan04 xband receiver moved to new position
(close to 310)
- 14jan04 610 receiver moved to new position
- 14jan04 laser ranging pc replaced with a new
- 14jan04 turret floor weighed. did not
include lbn or 6to8Ghz receiver.
- 14jan04 turret floor framing reinforced
(part that holds the rollers).
- 14jan04-20jan04 . 327 receiver was removed and then
re-installed when the turret floor was weighed.The cables from
the dewar were accidently left disconnected. They were
re-installed 21jan04 2am ast.
- 21jan04 PolB lbw has been having
glitches every 2.8 seconds. They are negative going in the
bandpass. They are coming from the dewar monitoring circuitry
affecting the bias voltage when the 70K dewar temp is read. The
70K dewar temp read was disable 21jan04 12:00 ast so the
glitches should no longer be present.
- 26jan04 leaf springs on dome changed out for
stronger springs.(date may be off by a week)..
- 27jan04 laser ranging pc moved to different
network switch. Archive file moved to observer.
- 28jan04 downstairs IF 720 to 780 Mhz filters
- 09feb04-08mar04 . vertex monitoring, tiedown monitoring were
not logging to disc.
- 10feb04 6 to 8 Ghz receiver installed on
- 10feb04 rfpwr3 volts to watts conversion
math9.c 723.c sband transmitter recalibrated.
- 11feb04 horn
(1st) all horns by lynn baker.
- 13feb04 cbh horn moved for focus correction.
- 13feb04 shigh horn moved down .8 inches to
bring into focus.
- 13feb04 shi new cal diodes installed.
- 14feb04 horn
survey after turret positions updated, cbh moved.
- 29feb04 updated cband high model offsets
using feb04 calib runs.
- 01mar04 replaced distomat 5 (this was the
last old distomat).
- 15mar04 xb rcvr reinstalled. had been on
hill. new cals, rfamps. fixed some loose cables in dewar.
- 16mar04 1320 hipass filter was installed with polA,polB
reversed. It was corrected today.
- 05apr04 xb receiver reinstalled. came down
29mar04 to work on dewar temp monitor circuit.
- 08apr04 turret floor reenforcing for alfa,
kevlar cables installed,tensioned
- 09apr04 resurveyed
tertiary into position using theodolite (alfa not yet in).
- 19apr04 (about). 3 new compressors raised to
dome for alfa. placed in new forward position on top of dome.
about 1000 lbs.
- 21apr04 alfa installed
- 23apr04 lbw
polB calval jumps.
- 28apr04 horn
survey, turret position updated for alfa.
- 29apr04 horns moved: xb,cbh,cb,sbh,sbw(x).
sbn,sbtx,lbw,alfa not moved.
- 01may04 tertiary
survey. updated tertiary origin.
- 02may04 horn
after receivers moved. Lynn had been using an "old"
secondary reference set for the last couple of surveys. Today he
switched back to the most recent. This caused a jump in y of
about .4 inches. The horn positioning above was done with the
- 05may04 low pass filters installed in alfa
- 06may04 alfa cover breaks on turret encoder
mount. lost encoder position. encoder remounted with a 85.47 deg
offset. inserted this in the little star program. tracked two
sources to get the turret encoder position close.
- 07may04 lbw had 56 asec error in za.
Probably from when the tertiary was moved and we did not move
lbw. In lbw model, i subtracted -56 asecs from za
- 08may04 resurvey some horns to get correct
turret encoder offset.
- 14may04 sbw cal diodes replaced
- 14may04 sbw focus motion done.
- 17may04 pointing model 15 completed (after verify run).
- 26may04 About 2:30 ast.. on generator, trouble with
power (60hz->57 hz). systems rebooting many times.
- 27may04 station clock jumped forward one second. Saw the
jump 10 seconds after midnite. Probably occured at midnite.
Reset the clock back at 8:24:30 ast. May have been the leap sec
flag set in track clock from the power resets of the previous
- 31may04 cband cal polA jumped . Tsys polA
went from 31 to about 48 K (jumped back 04jul04)
- 04jul04 cband polA jumped back. TsysPolA
went about 48 to 31. This occured after the dewar was warmed up
and then cooled back down. Problem must have been in the dewar.
Note.. continued jumping till 23sep04 when broken cable polA was
- 12jul04 motion failure azimuth axis. Caused
by loose/broken coupling on azimuth encoder. Replaced the
encoder and coupling. This is the encoder that is used for
positioning the azimuth arm (the encoder on the ch side is the
- 12aug04 ast 2:01:04 jpl hydrogen maser
stopped working. Off by 10 millisecs. 10:48:00 ast New maser
- 29aug04 galfa spectrometer installed.
- 08sep04 Rf to If board alfa pixel 0b
replaced. Had been an oscillation causing intermod near 1380
- 10sep04 new sbw gain curve,installed
10sep04. backdated to be valid starting15may04 (new cals
- 12sep04 lbw
cals measured 06aug04, installed 12sep04, backdated to
- 09oct04 cband polA Tsys had been jumping around
since may04. Beginning of oct04 dewar brought down and repaired.
there was a short in one of the wires for polA inside the dewar.
- 18nov04 sbn transmitter experiment.
Transmitter run but alfa was not covered.
- 23nov04 sbn 3db pad moved from after postamp
to before postamp. Put circulator after postamp. This was to fix
the out of band problems when a low side lo was used to mix 2380
to 750 Mhz. There had been lots of junk at around 800-1000 Mhz.
It was coming from the lousy match between the rf bandpass
filter and the postamp that followed it.
- 10dev04 updated epvchebs file
(/share/obs4/dt/dfiles/..) . old file was 1996 thru 2004. new
file is 2004 thru 2023. This holds the daily fits the the jpl
ephemeris that is used for the earth position info when
computing doppler shifts.
- 15dec04 18:45 az encoder 1 (dome side used for az pointing)
failed. Encoder was replaced.
- 27dec04 cbh: model offsets az,za incremented by
(-.28,-3.8) . data from x102 runs.
- 27dec04 60 mhz birdie. identified as sbwide postamp
chassis,cal injection board. Some unused hi speed chips were
- 30dec04: filters installed ch to get rid of tv station
- 17jan05 recalibrated load cells for tiedowns. replaced
filter td4-ldcell1 (top)
- 20jan05 azimuth would only go toward - az while in prgtrack
mode. preset, rate worked ok. Replaced lcu,plc. Started to work
ok. Found that the azencoder offset had been set to 1 degree
(since at least 06jun04). With the offset now at 1, the az again
trys to go to minus azimuth (exclusively). I reset the offset to
0 and then it worked ok. I also checked and on 19jan05, az=270,
with offset=1. the azencoder was the same as
20jan05,az=270, offset=0. So the software in the old lcu or plc
was ignoring the offset. The newer version is not? Another guess
is that the program track code in the plc is using the encoder
position before the offset is added for the pid loop. I saw that
a positive offset caused the az to only move to neg az, while a
negative offset would cause the az to only move to positive az
(for pt mode only).
- 27jan05: filter installed 430 Mhz system to get rid of tv
- 1feb05 : Noise injection boards for receivers.. all replaced
to get rid of 60Mhz birdie.
- 23feb05: (date not certain). lbw filter bank. 1370 hipass polB
hand a misthreaded n/sma connector, replaced. Also 1360-1470
bandpss pola had a loosed sma connector .. tightened it.
- 03mar05: lbw filterbank 1360-1470 filters. switched filters
polA->polB polB->polA to see if instabilities around 1390
change or the power dropouts.
- 04feb05: dome cam limit switches replaced with strikers and
regular switches when the cam switch broke.
- 25feb05: td 4 jack breaks. looks busted ball bearing.
replaced jack 28feb05. reset encoder, limit switches.
- 16mar05: model16 carriage house installed.
- 17mar05: td12 jack replaced (maintenance).
- 24mar05: alfa Rf to If converter pixel 0b replaced to get rid
of wandering birdie around 1430 Mhz.
- 29mar05: alfa rf to if convert pix0b spare. had been installed
without cover over chips. installed cover today.
- 05apr05: alfa pixel 2a,b 10db amp added before the mixer
upstairs to test oh performance.
- 08apr05: 10db amp installed in front of 1st mixer/filterbank
for all alfa pixels.
- 12apr05: 6db pads installed before the new 10db amps in alfa.
- 23apr05: wapp pulsar data archive started.
- 25apr05: alfa filter bank modified. 10db amps are now only on
the 1665 filters. the wideband filters are back to the way they
were before any of the filter bank changes.
- ??may05 : fans added to polA distribution amp downstairs.
(this may have caused the galfa cable switch).
- 25may05: cband receiver reinstalled. New amps installed, new
- 31may05 galfa pixel 4a and 6a (count from 0) were
switched at the output of the distribution amps downstairs. They
have been switched since at least 3may05. They were not switched
on 12dec04. At 11am ast this problem was corrected.
- 01jun05: lbw warmed up dewar. found loose/broken cable polA
probe. also found that cal cable polB in dewar was cracked.
(need to verify that these were the correctpols since the
banging caused polB to jump the most, not polA
- 09jun05: wapp4 has its kernel upgraded from 2.4.18 to 2.4.31.
- 10jun05: switches on dome wheels checking for failure enabled.
lights in control room. don't know if they are connected to the
emg stop circuit or not.
- 15jul05: wapps 1,2,3 have there kernel upgraded from 2.4.15?
- 21jun05: lbw installed new cal values. backdated to 01jun05.
(bug in d1->polB hi.. fixed 09aug05).
- 21jul05: diode 1 for lbw died. 22jul05 Switched the software
so no matter what people ask for they'll get diode 2. Will take
about 8 weeks to get some new diodes.
- 30jul05: installed new cal values for cband (amps/diodes
installed 25may05). The values for hcorcal polB look a little
- 03-04aug05: installed beams in feed tower to brace some beams
that are deflecting (probably caused by the kevlar cables).
- 08aug05: cband cals now correct (fixed bug in hcorcal/hxcal
- 09aug05: lbw cals .. fixed bug d1->polB high (first
- 18aug05: sbw cal cables changed. This modified the cal value
injected into the dewar
- 29sep05: updated cal values for 430ch (measured 27sep05).
- 30sep05: 0am till about 8:30 am (ast) station clock behind
real time by 1 second. leap second inadvertantly added to the
station clock. Pointing off by 1 sec of time (pointing at a
lower ra) .
- 11nov05: azimuth encoder problems enc1 and 2. on 14nov05 lost
encoder 1. realigned pointing with the index. 16nov05 replaced
- 23nov05: lbw new cal diodes installed.
- 28nov05: lbw new cal diode values installed.
- xxdec05: 1440 filter bank installed in alfa. xx gt 1
- 04dec05: lbw polB cal value jumped 20%. jump started on
04dec05, done by 05dec05.
- 23dec05: 430 yagi moved down .41 meters. Now 3 meters below
- 10jan06: air conditioner in rfi shack changed (old one was not
- 11jan06: put fan in distomat 6 (i could be off by a few days
on the date).
- 26jan06: 3rd ac installed in dome. 2 compressors are outside
at entry to dome . 1 compressor still on service platform. The
old ac is switched off. no longer needed.
- 23feb06: lbw cal values remeasured. installed on 28feb06 and
backdated to 04dec05 when polB cal jumped.
- 03apr06: aeron cpu, and 4 wapps. Removed cpu intensive jobs
being run every mornging at 4am by cron.daily (slocate.cron,
etc). Should help ntp time resets.
- 10apr06: sbn cals. Installed new values. backdated to 14oct03.
Need to reprocess old calibration data.
- 23may06: 430 yagi, found broken capacitor for 1 of the driving
elements. Ganesh remeasured the cal values (they were way off).
- 25may06: cband cal polA sma connector was loose. Tsys had been
wandering because of this.
- 25may06: 430 dome cal driver upstairs. Power supply had
failed. replaced it.
- 16jun06: lbw dewar warmed up and opened. found loose sma feed
thru in dewar for polB cal. Probably causing the drifts in polB
- 02aug06: new az2 encoder installed. create 614 khz comb in
- 28aug06: installed
for cband (measured back in 06jul06).
- 31aug06: installed
for lbw (measured on 06jul06). The cal change occurred on
16jun06 when the dewar was warmed up and a loose cal cable was
- 29sep06: removed final skybolt ap (ap#2) from the da vme
crate. Speeds up the reboot a lot. The ap was probably still
good (although it was no longer used).
- 05oct06: new az2 encoder replaced with old az encoder
(around1pm) to get rid of 614 khz comb in 327/430. It has been
there since 02aug06 (more
- 14oct06: az 1 encoder reading failed. Telescope switched to
use az2enc on ch side. This created a pointing error.
- 18oct06: 18:00. replaced az 1 encoder. realigned with index
pointer. Still bad. Found encoder board port 1 bad. Moved az1
encoder to encoder port 2 to get rid of pointing error
introduced by 14oct06 change.
- 26oct06: put repaired encoder board back in . Az 1 encoder now
working.not longer plugged into port 2. Got rid of az encoder
offset in lcu display (was 1 degree).
- 27oct06: foundloose
cable 750 mixer chassis polb. It was probably
causing jumps in polB iflo.
- 30oct06: replaced if attenuator polB upstairs. Failed 28oct06.
Was giving 22 db atten all the time. May have been causing
problems with polB if it was intermittantly failing before this.
- 14nov06: az encoder 1 failed. Had to replace it. Required a
realignment with the index.
- 14nov06: tiedown 12 loadcell 1 replaced with spare. It had
been about 25 ohms off at the loadcell.
- 27nov06: tertiary huffman box mounted on outside porch of
turret room brought down. weight:150 lbs.
- 28nov06: 3 cables removed from the catwalk tray: old sband
control cable,old telephone cable, and unused worn power cable.
Total weight removed: 1643 lbs.
- 5dec06: 05dec06 to 07dec06 removed 2115 lbs. Mainly from
bottom center of azimuth arm. Will be brought back up after
painting project. 245 lbs removed from the feed tower.
- 11dec06: updated
. Also updated the models for 1az
- 15feb07: crval1 bug in fits header fixed. Bug could have
header freq off by many Mhz. See cima page, cima fits
- 07mar07: bad
wapp4 brd1 polB (alfa beam6B). Fixed on 07mar07. It had
been bad during feb07.
- 31mar07: laser shack removed from corner 4 (about 1600 lbs) .
- 03apr07: lead from corner 4 removed. (5300 lbs) .
- 03apr07: lbw brought back up to the dome. screws omt tightened
hopefully getting rid of the 1400Mhz resonance.
- 04apr07: greg dome local disable cable breaks in dome cable
drape then bothazimuth
encoders reset .
- 13apr07: xband receiver brought down for maintenance.
- 14apr07: weather station fails. brought down for maintenance.
- 17apr07: alfa lowered from dome to electronics lab for
- 18apr07: vertical sides for painting containment installed
- 19apr07: vertical sides for painting containment installed
- 20apr07: weather station weather vane removed. Was in the way
of the painting. Will reposition it to another spot.
- 25apr07: stopped observing .. start of painting shutdown.
- 30apr07: stopped monitoring dewar temperatures (since no
- 04may07: looks like corner 12 containment sealed. Temp sensor
(inside containment > 110 degrees)
- 03dec07: start recommissioning telescope after painting
- 09dec07: moved nuts on td 4 and 8 to equalize td extension.
- 10dec07: alfa lifted and reinstalled
- 13dec07: started pointing model (data thru 16de07). 5
- 16dec07: finished pointing model
- 16dec07:evening started model verify.
- 04jan08: lifted 6th compressor to dome
- 09jan08: installed waveguide cable car landing to center
bearing (about 5000 lbs).
- 14jan08: lbw pointing. updated after verify run. offsets were
4.44 az, -2.68za asecs
- 05feb08:readjusted the pots for azimuth motor amplifiers so
they are at standard setting.
- 16apr08: increased alfa beam 0b gain by 3db at downstairs if
- 16apr08: 500 lbs placed on ch end
of azimuth arm . They put the work platform back.
- 17apr08: found loose cable downstairs synth1 used for mixer
lo's. At 1000Mhz output 10dbm at 1010 output 2dbm. Cable was
loose and the pin was short. Also found synth3 cable loose (but
pin was ok length).
- 13jun08: new cooled 327 receiver installed with new cal values
- 08jun08: lbw cryo
failures. About 4 warmups.
- 21jun08: updated
model using 1az,3az fits to the pointing data.
- 01jul08: subtracted 4asecs from xband za constant term for
bania. no other models were changed.
- 15aug08: 800 Mhz rfi feed installed in place of 610 feed.
- 29aug08: lbw cals measured before warmup. Installed backdated
- 05sep08: lbw warmed up, bias voltage bulkhead cable replaced.
- 19sep08: finished modifying alfa bias voltages for amp 3 for
long recovery times.
- 10oct08,20oct08: laser ranging reflectors replaced.
- 08jan09: td12 jack replaced.
- 01jul09: oplog az encoder off by 1 tooth??
- 02jul09: az 1 encoder replaced
- 15sep09: alfa down to be worked on
- 19sep09: td12 load cell broke. replaced 06oct09
- 30oct09: alfa reinstalled
- 03feb10: Found broken beam
platform where td12 backstay ties to platform (left side
looking from north). Not sure when it broke.
- 24mar10: repair on broken beam completed. still limited to 3.5
to 12 deg za
- 14sep10 : started
repairing u2-u2l3 by adding c channels.
- xxsep10: lbw window blown out. brought down to lab and fixed.
- 30sep10: alfa thermally cycled.
- 19oct10: alfa motor replaced.
- 21oct10: recompiled agcProg, tdtrk, pntProg to have new
software limits: 1.06 to 16 za.Probably won't install until
- 01nov10: cband gortex horn
comes loose. tsys goes up. fixed 17nov10.
- 05feb11: finished reinforcing the telescope.
- 08feb11: changed za limits back to 19.69 (from 16).
- 14mar11: 800
pointing model offsets updated (after mount bolt
- 22jun11 : interim
sbc3,4 Tsys increase from +5volt pwrsup failure of
halfbandfilter 3,4 chassis.
- 22jun11: if2 rack polB +5 volts power supply failure.
- 26jul11: alfa brought down to lab to work on it.
- 06sep11: alfa reinstalled
- 27feb12: thru 04mar12: pointing model 18 data taken (with new
- 09mar12:alfa brought down to the lab.
- 20apr12: alfa reinstalled
- 05jun12: model18A installed. now using new td positions.]
- 12aug12: lbw new cal values installed.
- 16aug12: xband new cal values installed.
- 19sep12: cbh new cal values installed.
- 27sep12: cband: new cal values installed
- 19oct12: 430 dome jennings relay polB removed. Little power
was getting through the relay.
- 28oct12: new sbn receiver
- 31oct12: horn survey sbn , sbtx receivers
- 31oct12: lbw 1370 hipass filter switches were causing ripples.
- 02nov12: horn survey all other receivers.
- 09nov12: sbn circulator postamp chassis had been installed
- 18nov12: lbw pressure relief valve blew out. Tsys problems lbw
started after this.
- 20nov12: sbn new cal values installed
- 26nov12: sbn moved up in z direction to put it in focus.
- 04dec12: lbw dewar brought down to lab to work on it.
- 07dec12: lbw dewar reinstalled in dome.
- 01apr13: punta borinquen
carsr radar starts testing.
- 29apr13: punta borinquen carsr radar now blanking in ao
- xxjul13 : sometime in jul13, punta borinquen carsr radar
stoppped blanking in our direction.
- xxjul13: faa carsr radar being brought online.
- 12jul13: dome drag
link comes loose
- 25jul13: ao19 cooled
receiver tested on telescope.
- 01aug13: sbn first stage of
amps burned up when shuter bypassed.
- 05aug13: turnstile junction 430 receiver damaged when ao19
- 08aug13: faa carsr radar starts broadcasting.
- 17sep13: 430dome polA bandpass filter reinstalled. fixed
solder joints, teflon connecteors
- 04oct13: punta borinquen carsr radar blanking toward ao
- 28oct13: faa radar starts blanking in ao direction (more
- 13jan14: 6.4 mag earthquake
damages cable m8-4 (more
- 11mar14: earthquake repair finished.
- 28mar14: alfa down to lab
- 21may14: alfa reinstalled
- 25nov14-09nov14 - scramnet observer cpu not forwarding packets
to cor,da crates (more
- 11mar15: alfa down to the lab
- xxjul14 : sbn amp replacement.the old sbh hi amp that
was installed 01aug13 was replaced by a new chalmers amplifier.
polB in dewar, polA outside
- 26mar15: hf mesh first raised. not hangking at final height
- 28apr15: alfa back up to the dome.
- 16sep15: site power. 750KVA transformers. changed taps on
output (by 2) to increase site voltage.
- 10jan16: az 1 encoder fails. replaced.
- 16feb16: observing stops.. platform painting. alfa down to the
- 21apr16: platform painting done..
- 04may16: first
fringes 12m, 305m
- 13may16: alfa backup to platform. beam 0a has problems..
- 18may16: new visitor center exhibits are now on during the day
- 24may16: ch cals hardwired to 192,122 K.. attenuator was not
- 02oct16: az encoder 1 fails. replaced.
- 10jul17 : sband heat exchangers remove from platform
- 04aug17: new sband heat exchangers installed on
- 25aug17: td4 jack replaced (more
- 20sep17: hurricane maria. line feed falls into dish
- 27oct17: 1400 lbs ballast added
end of az arm to compensate for missing line feed.
caltype selector fails. always corcal. fixed 5dec17
- 01dec17: model19A
- 09dec17: commercial power returns to ao.
- 15apr18: 3 trolleys from the dome cable drape were
pulled off of their beam.
- ran through the night with the trolleys hanging upside
- one of dome motor cables was damaged.
- 26-30may18: arcing from site power
phase A. At the power pole by the airglow lab (.ps) (.pdf)
- 180731: sband synth (upstairs) fails. Also caused trouble
since 10Mhz for 1st LO synth getting 10Mhz from sband synth
output.Not sure how long it was bad. Pulsar observers had seen
freq jumps for many months??
- 180824 : 2nd lo synth #3 was 12db low in output power at 2160
Mhz. at 1000 Mhz it was ok.. Not sure how long this was bad.
- 181230 maser fails. 14:30 utc started to fail. 18:20utc out of
- 181231 rubidium clock replaces maser 19:05 UTC.
- 190116: dust collector removed from platform (2130 lbs)
- 191127: sbn switched the stubs on the sbn omt, to reverse
polA,polB on the output of the dewar (more info)
- now: rhcp comes out on dewar chb, and lhcp comes out on
dewar cha (the tx transmits rchp)
- there was always another
- 200228 td4 jack fails. left jacks in place till new one
up and arunning on 03mar20. pos had td12 at 10"
- 200229 if2 synths fail after power dip. fixed freq till
200301 19:30 (only x102 lost).
- 200603 if2 synth 4 fails after power dip. 16:45
03jun20 till 8:00 08jun20 freq fixed at 550Mhz.
- 200810 auxmain4North fails.
- 220918 hurricane fiona.
- 220927 station clock was rubidium . saw 3ms jump vs utc (gps
ant providing 1sec tick had come down after fiona. switched to